A vapor compression cycle using R134a as the
refrigerant operates with evaporating and condensing
temperatures of -5°C and 40°C. Calculate (i) Carnot cycle
compression work, (ii) excess work due to the ‘superheat horn’,
(iii) the loss of work due to throttling, and (iv) the loss in
refrigeration capacity due to throttling.
Answer #1

Jiven that 4scaț conside ing tenpeaten is. 313 R113 a is a Vapou31 сопрел ssion cycle 313-268 Cop Desined effech wok input 3ha-2b8、24kl he=106-19 kJ kg ng Conside h-28535 kg then s shl=山34975 t 097786(24H . 31-43.-hi 231.86 kJ conside 239.8 -106 1 リ the supeJ7 heat honn,. due to workSubトence 288-35- dy 3) due to Supen heot- 21.s42 xy Cg to the loss of wosik S30.381 -1729 1Sefigesan capai in los ニ2411 31-106-2. 143.44 kJ = 30.9(-106.2 less in sefnieionCopariy is 14344-13342